OpenStudy (anonymous):

If a snowball melts so that its surface area decreases at a rate of 10 cm2/min, find the rate at which the diameter decreases when the diameter is 8 cm.

OpenStudy (campbell_st):

so this is a related rates question so is the snowball a sphere...

OpenStudy (baru):

answered same ques ^

OpenStudy (anonymous):

Yes it is a sphere campbell_st

OpenStudy (campbell_st):

ok... so using the diameter in stead of the radius in the surface area of a sphere it becomes \[S = \pi \times (\frac{d}{2})^2 ~~or~~ S = \frac{ \pi \times d^2}{4}\] does that make sense...?

OpenStudy (anonymous):

Oh! thanks baru...It's not exactly the same but I'll see

OpenStudy (anonymous):

Shouldn't it be 4pi(d/2)^2 ? Campbell_st

OpenStudy (campbell_st):

oops forgot the 4 \[S = 4 \pi (\frac{d}{2})^2 ~~~or~~S = \frac{4\pi d^2}{4} \] so the surface area in terms of diameter is \[S = \pi d^2\] sorry about the error

OpenStudy (campbell_st):

so can you find the derivative od S with respect to d...?

OpenStudy (anonymous):

That will be 2pi*d right?

OpenStudy (campbell_st):

correct so you now know \[\frac{dS}{dD} = 2\pi d\] you also know \[\frac{dS}{dt} = 10 \] you could use -10... but everything is decreasing... so you need to use \[\frac{dD}{dt} = \frac{dD}{dS} \times \frac{dS}{dt}\] so find the reciprocal of \[\frac{dS}{dD} ~~then~~multiply~~by~~\frac{dS}{dt}\] to get the equation forthe rate of change in the diameter with respect to time... then just substitute d = 8 for the answer... hope it helps

OpenStudy (anonymous):

So then the final answer would be 5/8pi ?

OpenStudy (campbell_st):

That's what I got... and it should be cm/min

OpenStudy (anonymous):

Thank you so much Campbell!!! :-D

OpenStudy (campbell_st):

glad to help